A triangle has base 13 feet (ft) and an altitude of 7 ft. Find its area?
A) 22.75 B) 91 C) 20 D) 45.5

Answer by richard1234(7193)

A = (1/2)bh = (1/2)(13 ft)(7 ft) = 45.5 ft^2, choice D.